OpenStudy (anonymous):

Rewrite the following equation in slope-intercept form: x + y = 5 y = x – 5 y = x + 5 y = -x + 5 y = -x -5

OpenStudy (anonymous):

so solve for y in the equation x + y = 5 and you get y = -x + 5

OpenStudy (aroub):

rewrite in the form y=mx+b